Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Intern. region
Vertex Tracker
Drift chamber
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (July 12, 2005)

Attendees: Rainer Bartoldus, Al Eisner, Fang Fang, Larry Gladney, Selina Li, Brian Lindquist, Dong Su, Dan Walker.


  • The new TSF board X15 was swapped with X1.
  • There were two L3 node trims due to L3 tracking since last Tuesday.


  • Rainer and Dan will be working together to get the prescale put in config. database.
  • Dan is working on L3 track matching to clean up events for the turn on curves that he showed last week for fastmon.


  • As mentioned above, the TSF-X15 was swapped with X1 to see if the TSF damaged events in X15 would happen again. see HN posted.
  • Currently, the ZPD damaged data are not filtered out in the data stream. Not sure how we should deal with this yet.
  • Since the damage monitoring was installed at JAS, shifters are able to see that the ZPD and TSF damaged events showed up every once a while, but all of the occurances in the past week or so are only a few events per run. We have increased the threshold for paging the commissioner to 50 entries for damaged events since real problems would tend to persist and last longer. The TSF damaged events with out-of-frame errors. We don't understand the GLINK errors for ZPD that happened this week in one of the runs with only 3 damaged events in ZPD2. The short term instruction for DQM was updated. The commissioner will get paged when there are more than 50 entries in the damage monitoring plot, and email will be sent to babartrg if less than 50 entries so that we can keep track of this problem. It was suggested that a damage time history plot be shown to help monitoring this.
  • Al reminded us of the TSF raw data not being simulated in MC which was discussed long time ago. Originally, Nicolas was waiting for a decision to be made before proceeding. Rainer and Su Dong thought that the TSF segments are not in mini mainly because the data volume is substantial.


  • Su Dong is working on the fast control.


  • To avoid getting L3 node trim due to L3 tracking, we should probably restore the L3 tracking code to the one we used in Run4. We will make a decision in this week's L3 meeting on Friday.

Data Quality and Monitoring

  • Brian is working on time history plots for variables such as TSF segments and ZPD strips etc. within a run for fastmon.
  • We are checking the reprocessed Run2 and Run4 data, as well as Run5.

This page is created by Selina Li (; however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.